### v2

@@ 1. $(*_{3})$
[Why? As $n_{2}>n_{1}\geqslant n_{*}$ and so $\mathrm{rk}(f_{(\omega,n_{2})}(\langle X\rangle^{*}_{G}))=\mathrm{rk}(f_{(\omega,n_{1})(\langle X\rangle^{*}_{G})})$, recalling that $f_{(n_{2},n)}$ maps $f_{(\omega,n_{2})}(\langle X\rangle^{*}_{G})$ onto $f_{(\omega,n_{1})}(\langle X\rangle^{*}_{G})$, so the map $f_{(n_{2},n_{1})}\restriction f_{(\omega,n_{2})}(\langle X\rangle^{*}_{G})$ has trivial kernel and so it is $1$-to-$1$, and trivially it is also onto.]
@@ 1. $(*_{6})$
      Let $Y_{n}\subseteq T\cap\omega^{n}$ be finite and such that we have:
      $f_{(\omega,n)}(\langle X\rangle^{*}_{G})\subseteq G_{(n,Y_{n})}=\bigoplus\{K_{\nu}x_{\nu}:\nu\in Y_{n}\}.$
@@ 1. $(*_{8})$
    $(\mathrm{rk}(\pi_{(n,Y^{\star}_{n})}\circ f_{(\omega,n)}(\langle X\rangle^{*}_{G})):n\geqslant n_{*})$ is eventually constant, say for $n\geqslant n_{\star}$, and let $k_{\star}$ be this constant value.
@@ 2. $(\cdot_{2})$
      if $\nu\in Y^{\star}_{n}$ and $\nu\triangleleft\rho\in Y_{n_{2}}$, then $\rho\in Y^{\star}_{n_{2}}$.
@@ 1. $(*_{12})$
    For every $n<\omega$ and $\ell<k_{*}$, let:
    $y_{(\ell,n)}=\pi_{(n,Y^{\star}_{n})}\circ f_{(\omega,n)}(y_{\ell}),$
where $\pi_{(n,Y^{\star}_{n})}:G_{n}\rightarrow\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}$ is the obvious projection onto $\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}$.
@@ 1. $(*_{16})$
As in the beginning of the proof we can carry the induction. Now, as we are assuming that $S$ has no $\omega$-branch, clearly for some $n_{\dagger}>n_{+}$ we have that $\nu_{n_{\dagger}}\notin S$, but then necessarily $L(\nu_{n_{\dagger}})$ has be finite, i.e., $\nu_{n}\in Y^{\star}_{n}$, since by assumption $S$ is a subtree of $T$ and we defined, for $\eta\in T$, $L(\eta)=\{p_{\nu}:(\nu\triangleleft\eta\text{ or }\eta\trianglelefteqslant\nu)\text{ and }\nu\in S^{+}\}$.
@@
$\begin{array}[]{rcl}\pi_{(n_{\dagger},Y^{\star}_{n_{\dagger}})}\circ f_{(\omega,n_{\dagger})}(z)&=&\pi_{(n_{\dagger},Y^{\star}_{n_{\dagger}})}\circ f_{(\omega,n_{\dagger})}(\sum\{b_{\ell}y_{\ell}:\ell\leqslant k_{\star}\})\\ &=&\sum\{b^{n_{\dagger}}_{\ell}y_{(n_{\dagger},\ell)}:\ell\leqslant k_{\star}\}\\ &=&\sum\{b_{\ell}y_{(n_{\dagger},\ell)}:\ell\leqslant k_{\star}\}\\ &=&0\\ \end{array},$
@@ 1. $(*_{17})$
      $f^{\star}_{(\omega,n)}$ is an homomorphism from $G$ into $\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}\leqslant G_{n}$.
@@ 1. $(*_{18})$
    If $n\geqslant n_{*}$, then $g_{0}=f^{\star}_{(\omega,n)}\restriction\langle X\rangle^{*}_{G}$ is $1$-to-$1$ from $\langle X\rangle^{*}_{G}$ onto $\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}$.
@@ 1. $(*_{19})$
    $(*_{1})$ holds and so the proof is complete.
@@
4. Proofs of main theorems
@@
Let $\mathbb{P}_{*}$ be an infinite set of primes. Let $S$ be a well-founded (i.e., no $\omega$-branches) subtree of $\omega^{<\omega}$ such that for every $\eta\in S$ we have that $\{\nu\in S:\eta\triangleleft\nu\}$ is infinite and $0\notin\mathrm{ran}(\eta)$. We define a pair $(\omega^{<\omega},L_{\mathbb{P}_{*}})$ as in 3.10 by letting $T=\omega^{<\omega}$, $S\subseteq T$ the tree we just fixed and $(p_{\eta}:\eta\in S)$ be a list of $\mathbb{P}_{*}$ without repetitions.
@@
By 3.15, for $(\omega^{<\omega},L_{\mathbb{P}})$ as in 4.1, $G(\omega^{<\omega},L_{\mathbb{P}})$ (cf. 3.2) is $\aleph_{1}$-free.
We need the following definition from [11].
Definition 4.3.
For a category $\mathcal{C}$ by Pro-$\mathcal{C}$ we denote the category of $\omega$-inverse systems where the morphisms are given as sequences of onto maps $f_{n}:A_{\varphi(n)}\to B_{n}$ where $\varphi:\omega\rightarrow\omega$ is increasing, and all is commutative, that is, letting $A=(A_{n},\pi^{A}_{(n,m)}:m\leqslant n<\omega)$ and $B=(B_{n},\pi^{B}_{(n,m)}:m\leqslant n<\omega)$ in Pro-$\mathcal{C}$, for every $n\leqslant m<\omega$, we have that $f_{m}\circ\pi^{A}_{(\varphi(n),\varphi(m))}=\pi^{B}_{(n,m)}\circ f_{n}$.
@@
[…]
Proof.
Suppose that $\gamma:G\rightarrow H$ is a continous surjection (epimorphism). Recall that procountable group means that there is a nbhd basis of $e_{G}$ of normal open subgroups. So if $G$ is procountable, $D$ is discrete and $f:G\to D$ is a continuous epimorphism, then the kernel of $f$ being an open subgroup contains a subgroup from this nbhd basis. This yields the desired morphism of Pro-Groups from $A$ to $B$. ∎
Lemma 4.5.
Let $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ be infinite sets of primes such that $\mathbb{P}_{1}\cap\mathbb{P}_{2}$ is finite, and let $(\omega^{<\omega},L_{\mathbb{P}_{\ell}})$ be as in 4.1 and $G_{\ell}=G(\omega^{<\omega},L_{\mathbb{P}_{\ell}})$ (recall 3.2). Then there is no continuous surjection from $G_{1}$ onto $G_{2}$.
@@
We recall that by product group we mean a group of the form $\prod_{n<\omega}H_{n}$.
@@
[…]
@@
Let $A\subseteq\omega^{<\omega}$ be an infinite tree (in this case, crucially, possibly with leaves). For $\eta\in A$, let $\eta^{+1}=(\eta(n)+1:n<\omega)$ and let $A^{+1}=S=\{\eta^{+1}:\eta\in A\}$. Let $\bar{p}=\{p_{\eta}:\eta\in S\}$ and $T=\omega^{<\omega}$. And let $G(\omega^{<\omega},L_{A})$ be defined as in 3.10 for this choice of $S$ and $\bar{p}$.
@@
For $(\omega^{<\omega},L_{A})$ as in 4.6, letting $G_{A}=G(\omega^{<\omega},L_{A})$ we have:
@@
This is by 2.12 and 4.7, together with the observation that the map $A\mapsto\mathrm{inv}(\omega^{<\omega},L_{A})$ is Borel, when considering models whose domain is $\subseteq\omega$. ∎
@@ 5.1. Separating $\mathbb{Z}$-homogeneous and $\aleph_{1}$-free
In this section we prove Theorem 1.4, this result is of independent interest.
@@
    for every $X\subseteq K_{\mathbf{n}}$ of size $<\mathbf{n}$ we have that $\langle X\rangle^{*}_{K_{\mathbf{n}}}$ is free;
@@
    $\langle\bar{x}\rangle^{*}_{K_{\mathbf{n}}}$ is not free, in fact, more strongly, if $\mathbb{P}_{0}\subseteq\mathbb{P}_{*}$, then the group
    $K_{(\mathbf{n},\mathbb{P}_{0})}=\langle x_{\ell},y_{p}:\ell<\mathbf{n}\text{ and }p\in\mathbb{P}_{0}\rangle_{K_{\mathbf{n}}}\leqslant\langle\bar{x}\rangle^{*}_{K_{\mathbf{n}}}$
    is not free and disjoint from $\{y_{p}:p\in\mathbb{P}_{*}\setminus\mathbb{P}_{0}\}$.
@@
    for each $n<\omega$, there is $g_{n}:\omega\rightarrow\omega$ onto such that for $j<\omega$:
    $h_{(n,j)}:=f_{(n+1,n}\restriction G_{(n+1,j)}\in\mathrm{Hom}(G_{(n+1,j)},G_{(n,g_{n}(j))});$
@@
    $g_{n}\restriction\{2j+2:j<\omega\}$ is $1$-to-$1$ onto $\omega\setminus\{0\}$ and $h_{(n,2j+2)}(x_{(n+1,2j+2)})=x_{g_{n}(2j+2)}$.
@@ 5.2. For every $\eta\in[T]$, $\bigcap\{L(\eta\restriction n)\}$ is finite is not sufficient
Case 2. $n=m+1$ Now $(p_{\eta},Q_{\eta}:\eta\in 2^{m})$ are determined by the other conditions, but we have to find for $\eta\in 2^{m}$, the numbers $a_{\eta^{\frown}(0)},a_{\eta^{\frown}(1)}\in\mathbb{Z}\setminus\{0\}$ satisfying (d)-(f). To this extent, choose $a^{\prime}_{0},a^{\prime}_{1}\in\mathbb{Z}$ such that:
@@
Let $y_{n}=\sum\{a_{\eta}x_{\eta}:\eta\in 2^{n}\}$. Now, clearly $y_{n}\in G_{n}$. Furthermore, if $n<m<\omega$, then $f_{(m,n)}(y_{m})=y_{n}$. To see this it suffices to prove it for $n=m+1$. To this extent notice:
$\begin{array}[]{rcl}f_{(n+1,n)}(y_{n})&=&f_{(n+1,n)}(\sum\{a_{\eta}x_{\eta}:\eta\in 2^{n+1}\})\\ &=&\sum\{a_{\eta}f_{(n+1,n)}(x_{\eta}):\eta\in 2^{n+1}\}\\ &=&\sum\{a_{\eta}x_{\eta\restriction n}:\eta\in 2^{n+1}\}\\ &=&\sum\{\sum\{a_{\eta}:\nu\triangleleft\eta\in 2^{n+1}\}:\nu\in 2^{n}\}\\ &=&\sum\{(a_{\nu^{\frown}(1)}+a_{\nu^{\frown}(0)})x_{\nu}:\nu\in 2^{n}\},\\ \end{array}$

### v3

@@ 1. $(*_{3})$
[Why? As $n_{2}>n_{1}\geqslant n_{*}$ and so $\mathrm{rk}(f_{(\omega,n_{2})}(\langle X\rangle^{*}_{G}))=\mathrm{rk}(f_{(\omega,n_{1})(\langle X\rangle^{*}_{G})})$, recalling that $f_{(n_{2},n)}$ maps $f_{(\omega,n_{2})}(\langle X\rangle^{*}_{G})$ onto $f_{(\omega,n_{1})}(\langle X\rangle^{*}_{G})$, we have that the map $f_{(n_{2},n_{1})}\restriction f_{(\omega,n_{2})}(\langle X\rangle^{*}_{G})$ has trivial kernel and so it is 11-to-11, and trivially it is also onto.]
@@ 1. $(*_{6})$
      Let $Y_{n}\subseteq T\cap\omega^{n}$ be finite and such that we have: f_(ω, n)(⟨X ⟩^*_G) ⊆G_(n, Y_n) = ⨁{ K_νx_ν: ν∈Y_n}.
@@ 1. $(*_{8})$
    Recalling the definition of $\pi_{(n,Y^{\star}_{n})}$ from $(*_{1})$, we have that the sequence (rk(π_(n, Y^⋆_n) ∘f_(ω, n)(⟨X⟩^*_G)) : n ⩾n_*) is eventually constant, say for $n\geqslant n_{\star}$, and let $k_{\star}$ be this constant value.
@@ 2. $(\cdot_{2})$
      if $\nu\in Y^{\star}_{n_{1}}$ and $\nu\triangleleft\rho\in Y_{n_{2}}$, then $\rho\in Y^{\star}_{n_{2}}$ ($Y^{\star}_{n_{1}}$ was defined in $(*_{7})$).
[Why $(*_{9})(\cdot_{1})$? By $(*_{6})$(c).]
@@ 1. $(*_{12})$
    For every $n<\omega$ and $\ell<k_{*}$, let: y_(ℓ, n) = π_(n, Y^⋆_n) ∘f_(ω, n)(y_ℓ),
where $\pi_{(n,Y^{\star}_{n})}:G_{n}\rightarrow\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}$ is the obvious projection onto $\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}$ (as in $(*_{1})$).
@@ 1. $(*_{16})$
As in the beginning of the proof we can carry the induction. Now, as we are assuming that $S$ has no $\omega$-branch, clearly for some $n_{\dagger}>n_{+}$ we have that $\nu_{n_{\dagger}}\notin S$, but then necessarily $L(\nu_{n_{\dagger}})$ has to be finite, i.e., $\nu_{n}\in Y^{\star}_{n_{\dagger}}$, since by assumption $S$ is a subtree of $T$ and we defined, for $\eta\in T$, $L(\eta)=\{p_{\nu}:(\nu\triangleleft\eta\text{ or }\eta\trianglelefteqslant\nu)\text{ and }\nu\in S^{+}\}$.
@@
[…]
@@ 1. $(*_{17})$
      $f^{\star}_{(\omega,n)}$ is a homomorphism from $G$ into $\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}\leqslant G_{n}$.
@@ 1. $(*_{18})$
    If $n\geqslant n_{*}$, then $g_{0}=f^{\star}_{(\omega,n)}\restriction\langle X\rangle^{*}_{G}$ is 11-to-11 from $\langle X\rangle^{*}_{G}$ onto $\langle Y^{\star}_{n}\rangle^{*}_{G_{n}}$.
@@ 1. $(*_{19})$
    The conditions from $(*_{1})$ holds and so the proof is complete.
@@
4. Proof of Main Theorems
@@
Let $S$ be the following well-founded tree. For $n<\omega$, let $T_{n}$ be a copy of $\omega^{<n}$ such that for $n<m<\omega$ we have that $T_{n}\cap T_{m}=\emptyset$. Let $S$ be the tree obtained adding a minimum $x_{*}$ to $\bigcup\{T_{n}:n<\omega\}$, so $x_{*}$ is the root of the tree $S$. Easily, $S$ can be realized as a subtree of $\omega^{<\omega}$ satisfying condition 3.9(2), and $S$ is well-founded. Fix an enumeration of $S$ of order type $\omega$, say $(\eta_{i}:i<\omega)$. Let $\mathbb{P}_{*}$ be an infinite set of primes, enumerate $\mathbb{P}_{*}$ in increasing order as $(p^{\mathbb{P}_{*}}_{i}:i<\omega)$ and let $p_{\eta_{i}}=p^{\mathbb{P}_{*}}_{i}$. Then $S$ and $\bar{p}=(p_{\eta_{i}}:i<\omega)$ are as in Definition 3.9. W.l.o.g. we assume that if $\eta$ is not the root of the tree, then $\eta(0)=n$ if and only if $\eta\in T_{n}$.
@@
By 3.14, for $(\omega^{<\omega},L_{\mathbb{P}_{*}})$ as in 4.1, $G(\omega^{<\omega},L_{\mathbb{P}_{*}})$ is $\aleph_{1}$-free.
@@
Let $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ be infinite sets of primes such that $\mathbb{P}_{1}\cap\mathbb{P}_{2}$ is finite, and let $(\omega^{<\omega},L_{\mathbb{P}_{\ell}})$ be as in 4.1 and $G_{\ell}=G(\omega^{<\omega},L_{\mathbb{P}_{\ell}})$ (recall 3.2). Then there is no topological isomorphism from $G_{1}$ onto $G_{2}$.
@@
  1. (a)
    $\phi:\omega\rightarrow\omega$ is increasing;
  2. (b)
    $\psi:\omega\rightarrow\omega$ is increasing;
  3. (c)
    $f_{n}:A_{\phi(n)}\rightarrow B_{n}$;
  4. (d)
    $g_{n}:B_{\psi(n)}\rightarrow A_{n}$;
  5. (e)
    for every $n\leqslant k<\omega$, the following diagram is commutative:
    $A_{\phi(k)}$$B_{k}$$A_{\phi(n)}$$B_{n}$$\scriptstyle{f_{k}}$$\scriptstyle{t_{(\phi(k),\phi(n))}}$$\scriptstyle{s_{(k,n)}}$$\scriptstyle{f_{n}}$
  6. (f)
    for every $n\leqslant k<\omega$, the following diagram is commutative:
    $B_{\psi(k)}$$A_{k}$$B_{\phi(n)}$$A_{n}$$\scriptstyle{g_{k}}$$\scriptstyle{s_{(\psi(k),\psi(n))}}$$\scriptstyle{t_{(k,n)}}$$\scriptstyle{g_{n}}$
  7. (g)
    for every $n<\omega$, $f_{n}\circ g_{\phi(n)}:B_{\psi(\phi(n))}\rightarrow B_{n}$ is equal to $s_{(\psi(\phi(n)),n)}$;
  8. (h)
    for every $n<\omega$, $g_{n}\circ f_{\psi(n)}:A_{\phi(\psi(n))}\rightarrow A_{n}$ is equal to $t_{(\phi(\psi(n)),n)}$.
  1. $(*_{1})$
    If $\eta\in\omega^{\phi(k)+1}$ is not maximal in $S$, then $f_{k}(x_{\eta})=0_{B_{k}}$.
[Why $(*_{1})$? By our choice $\eta$ is not maximal in $S$, hence we have that:
[…]
but since $\mathbb{P}_{1}\cap\mathbb{P}_{2}$ is finite, necessarily $f_{k}(x_{\eta})=0_{B_{k}}$, by 3.7.]
Let $n=\phi(\psi(0))$ and let $\eta\in\omega^{n+1}$ be such that $\eta$ is not maximal in the tree $S$. Then by $(*_{1})$ we have that $f_{\psi(n)}(x_{\eta})=0_{B_{\psi(0)}}$ and so necessarily $g_{0}\circ f_{\psi(0)}(x_{\eta})=0_{A_{0}}$, but this contradicts item (h) above, in fact, since $g_{n}\circ f_{\psi(n)}:A_{\phi(\psi(0))}\rightarrow A_{0}$ is equal to $t_{(\phi(\psi(0)),0)}$ we also have that $g_{0}\circ f_{\psi(0)}(x_{\eta})=x_{\emptyset}$ and $x_{\emptyset}\neq 0_{A_{0}}$, a contradiction. ∎
@@ 1. $(*_{1})$
Relying on 4.3, it suffices to take a collection $\{\mathbb{P}_{\alpha}:\alpha<2^{\aleph_{0}}\}$ of almost disjoint subsets of the set of prime numbers, where $A,B$ are almost disjoint if $A\cap B$ is finite. ∎
@@
Let $A\subseteq\omega^{<\omega}$ be an infinite tree (in this case, crucially, possibly with leaves). For $\eta\in A$, let $\eta^{+1}=(\eta(n)+1:n<\omega)$ and let $A^{+1}=S=\{\eta^{+1}:\eta\in A\}$. Let $\bar{p}=\{p_{\eta}:\eta\in S\}$ and $T=\omega^{<\omega}$. And let $G(\omega^{<\omega},L_{A})$ be defined as in 3.9 for this choice of $S$ and $\bar{p}$.
@@
For $(\omega^{<\omega},L_{A})$ as in 4.4, letting $G_{A}=G(\omega^{<\omega},L_{A})$ we have:
@@
First of all, recall that, as well-known, well-founded trees form a complete co-analytic set. This is by 2.13 and 4.5, together with the observation that the map $A\mapsto\mathrm{inv}(\omega^{<\omega},L_{A})$ is Borel, when considering models whose domain is $\subseteq\omega$. ∎
@@ 5.1. Separating $\mathbb{Z}$-homogeneous and $\aleph_{1}$-free
In this section we prove Theorem 1.4. This result is of independent interest.
@@
    for every $X\subseteq K_{\mathbf{n}}$ of size $\mathbf{n}$ we have that $\langle X\rangle^{*}_{K_{\mathbf{n}}}$ is free;
@@
    $\langle\bar{x}\rangle^{*}_{K_{\mathbf{n}}}$ is not free, in fact, more strongly, if $\mathbb{P}_{0}\subseteq\mathbb{P}_{*}$, then the group K_(n, P_0) = ⟨x_ℓ, y_p : ℓ¡ n and p ∈P_0 ⟩_K_n ⩽⟨¯x ⟩^*_K_n is not free and disjoint from $\{y_{p}:p\in\mathbb{P}_{*}\setminus\mathbb{P}_{0}\}$.
@@
    for each $n<\omega$, there is $g_{n}:\omega\rightarrow\omega$ onto such that for $j<\omega$: h_(n, j) := f_(n+1, n ↾G_(n+1, j) ∈Hom(G_(n+1, j), G_(n, g_n(j)));
@@
    $g_{n}\restriction\{2j+2:j<\omega\}$ is 11-to-11 onto $\omega\setminus\{0\}$ and $h_{(n,2j+2)}(x_{(n+1,2j+2)})=x_{g_{n}(2j+2)}$.
@@ 5.2. For every $\eta\in[T]$, $\bigcap\{L(\eta\restriction n)\}$ is finite is not sufficient
Case 2. $n=m+1$ Now $(p_{\eta},Q_{\eta}:\eta\in 2^{m})$ are determined by the other conditions, but we have to find for $\eta\in 2^{m}$, the numbers $a_{\eta^{\frown}(0)},a_{\eta^{\frown}(1)}\in\mathbb{Z}\setminus\{0\}$ satisfying (d)-(f). To this end, choose $a^{\prime}_{0},a^{\prime}_{1}\in\mathbb{Z}$ such that:
@@
Let $y_{n}=\sum\{a_{\eta}x_{\eta}:\eta\in 2^{n}\}$. Now, clearly $y_{n}\in G_{n}$. Furthermore, if $n<m<\omega$, then $f_{(m,n)}(y_{m})=y_{n}$. To see this it suffices to prove it for $n=m+1$. For this purpose, notice:
[…]